OpenStudy (anonymous):

The Volume of a cuboid is 84cm^3 with the height of 7cm. The least possible length of the diagonal of the base of cuboid is

OpenStudy (anonymous):

(A)\[\sqrt{23}\] (B)\[\sqrt{24}\] (C)\[\sqrt{25}\] (D)\[\sqrt{144}\]

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

ia min class 9 i don't understand this

OpenStudy (raden):

assume the base of cuboid is a square with length is x, so we have x^2*7=84 x^2 = 84/7=12 x=sqrt(12) now, to find the diagonal of base, u can use the formula: sqrt(2) * side base = sqrt(2) * sqrt(12) = ....

OpenStudy (anonymous):

\[\sqrt{24}\]

OpenStudy (raden):

right

OpenStudy (anonymous):

just 1 question didn't understand ur now, to find the diagonal of base, u can use the formula: sqrt(2) * side base part

OpenStudy (anonymous):

i think i understood is this diagonal of a square formulae?

OpenStudy (raden):